OpenStudy (anonymous):

c^2-4c=0, what is the solutions to this equation? 0,-4 0,4(square rooted) 0,4 1,-4(square rooted)

OpenStudy (anonymous):

are*

OpenStudy (anonymous):

0,4

OpenStudy (anonymous):

How is it that :3

OpenStudy (anonymous):

\[\huge c^2-4c=0 \rightarrow c(c-4)=0\] \[\huge c= 0 ---or ---(c-4)=0\] c=0 or c= 4 Hence 0, 4 are the required answers

OpenStudy (anonymous):

Gotcha, thanks. I'm unsure because I didn't know how to do it with a binomial.